The relationship between angular frequency and wavenumber for a deep water wave is:

?=gk????

where g is the acceleration due to gravity. If the phase velocity of a particular wave is 9.5 m/s what is the group velocity of the wave?
Answer #1

Vp*Vg=c^2

Vg=c^2/Vp=9*10^16/9.5=9.47*10^15 m/s

In non dispersive frame

Vp=Vg

Vg=9.5m/s